# Long-context LLMs Struggle with Long In-context Learning

### v2

@@ 4 Exploratory Experiment
Inspired by the Lost in the Middle phenomenon  Liu et al. (2023), we take analysis experiments to explore whether the position distribution of the instances will make a difference in the performance for long in-context learning with extreme-label classification tasks.
@@ 4.1 Scattered Distribution
In our investigation, we conducted pilot experiments on TacRED, a medium-complexity dataset, with each label type demonstrated three times, resulting in a total of 123 distinct instances (calculated as $41\times 3$). Within these experiments, instances bearing the same labels are distributed randomly to form a scattered configuration. For each instance, we track its relative position within the prompt alongside its corresponding label, thereafter computing the accuracy for each label class. As illustrated in the first row of Figure 4, the visualization delineates the accuracy of each label, aligned with its position within the prompt, where diverse colors symbolize various label types. In scenarios where class instances are scattered, certain models, such as InternLM2-7B-base, demonstrate acceptable performances—approximately 60% accuracy merely on specific labels, as highlighted by a red circle in Figure 4, regardless of the instance placements. Conversely, other models, like ChatGLM3-6B-32K, exhibit robust performance across a broad spectrum of labels. Remarkably, the GPT4-turbo model consistently surpasses an 80% accuracy threshold for the majority of label types, with only a minimal count of exceptions.
@@ 4.2 Grouped Distribution
To facilitate a clear comparison between scattered and grouped distributions, we organize instances of the same class to be adjacent within the demonstration prompts. The impact of this reorganization on model performance, both pre and post-grouping, is presented in Table 7. A pronounced trend emerges, highlighting a general decline in performance across most models after grouping instances by class. Notably, models such as Mistral-7B-v0.2-base and InternLM2-7B-base exhibit significant performance drops, underscoring a pronounced sensitivity to instance grouping. In an effort to delve deeper into this phenomenon, we visualize the accuracy of grouped labels in relation to their positions within the prompt, as illustrated in Figure 4. This visualization reveals that instances of the same class, denoted by dots of the same color, are positioned nearby. It became evident that some models, like InternLM2-7B-base, demonstrate high sensitivity to the distribution of instances, only handling instances with labels positioned at the end of the prompt. Conversely, other open-source models such as ChatGLM3-6B-32K, with a modest 3.3% drop in accuracy, proved to be more resilient to changes in instance positioning, maintaining high performance across varied positions. Surprisingly, even the GPT4-turbo is not immune to the challenges posed by grouped distributions, experiencing a notable decline in performance by 20.3%. This observed decrease in performance is consistent across models, unaffected by the specific positions of the labels within the prompt.
Model Param Support Scatter Grouped $\Delta$
Context Tokens 10K
Gemma-7B-base 7B 8K 0 0 0
LLaMA-2-7B-32K 7B 32K 0.4 3.0 +2.6
ChatGLM3-6B-32K 6B 32K 38.9 35.6 -3.3
Qwen-1.5-7B-base 7B 32K 45.2 33.0 -12.2
Mistral-7B-v0.2-base 7B 32K 51.6 5.1 -46.5
LLaMA-2-7B-LongLora 7B 100K 0 0 0
Yi-6B-200K 6B 200K 8.0 0 -8
InternLM2-7B-base 7B 200K 15.5 4.8 -9.7
Long-LLaMA-code-7B 7B 256K 4.1 0 -4.1
RWKV-5-World 7B 4K 1.0 3.6 +2.6
Mamba-2.8B 2.8B 2K 0 0 0
GPT4-turbo N/A 128K 79.5 59.2 -20.3
@@ 5 Conclusion
In summary, our research explores the capability of large language models on long in-context learning tasks, particularly in extreme-label classification scenarios. We curate a dataset LongICLBench consisting of long in-context learning tasks with different difficulty levels with respect to the context length. Through our study, we have discovered that while LLMs show promising performance on inputs up to 20K tokens, their ability to process and understand longer sequences significantly decreases. Our exploratory experiments further highlight the impact of the distribution of examples within prompts on model performance. We hope LongICLBench and our findings contribute to the ongoing efforts to enhance LLMs’ understanding of long contexts.
@@ A.2 Prompting Template
{comment: ”…comment…”
emotion category: ”…emotion…”
@@
{service query: ”…service…”
intent category: ”…intent…”
@@
{sentence: ”…sentence…
the subject is ”…subject…”
the object is ”…object…”
the relation between the two entities is: ”…relation…”
@@
{”…entity…”: ”…entity type…”
@@
”…dialogue…”
The list of entity pairs are ”…(subject1, object1), (subject2, object2), etc…
The ”…number of pairs…” respective relations between each entity pair are: ”…relation, relation2, etc…
@@
{”…sentence1…” ( ) ”…sentence2…”
the conjunction word in ( ) is ”…conjunction…”

### v3

@@ 4 Exploratory Experiment
Inspired by the Lost in the Middle phenomenon (Liu et al., 2023), we take analysis experiments to explore whether the position distribution of the instances will make a difference in the performance for long in-context learning with extreme-label classification tasks.
@@ 4.1 Scattered Distribution
In our investigation, we conduct pilot experiments on TacRED, a medium-complexity dataset, with each label type demonstrated three times, resulting in a total of 123 distinct instances (calculated as $41\times 3$). Within these experiments, instances bearing the same labels are distributed randomly to form a scattered configuration. For each instance, we track its relative position within the prompt alongside its corresponding label, thereafter computing the accuracy for each label class. As illustrated in the first row of Figure 4, the visualization delineates the accuracy of each label, aligned with its position within the prompt, where diverse colors symbolize various label types. In scenarios where class instances are scattered, certain models, such as InternLM2-7B-base, demonstrate acceptable performances—approximately 60% accuracy merely on specific labels, as highlighted by a red circle in Figure 4, regardless of the instance placements. Conversely, other models, like ChatGLM3-6B-32K, exhibit robust performance across a broad spectrum of labels. Remarkably, the GPT4-turbo model consistently surpasses an 80% accuracy threshold for the majority of label types, with only a minimal count of exceptions.
@@ 4.2 Grouped Distribution
To facilitate a clear comparison between scattered and grouped distributions, we organize instances of the same class to be adjacent within the demonstration prompts. The impact of this reorganization on model performance, both pre and post-grouping, is presented in subsection A.3. It is easy to observe that there is a general decline in performance across most models after grouping instances by class. Notably, models such as Mistral and InternLM2 exhibit significant performance drops, underscoring a pronounced sensitivity to instance grouping. In an effort to delve deeper into this phenomenon, we visualize the accuracy of grouped labels in relation to their positions within the prompt, as illustrated in Figure 4. This visualization reveals that instances of the same class, denoted by dots of the same color, are positioned nearby. It became evident that some models, like InternLM2 or Mistral shown in subsection A.3, demonstrate high sensitivity to the distribution of instances, only handling instances with labels positioned at the end of the prompt. Conversely, other open-source models such as ChatGLM3-6B-32K, with a modest 3.3% drop in accuracy, proved to be more resilient to changes in instance positioning. Surprisingly, even the GPT4-turbo and Gemini1.5-Pro are not immune to the challenges posed by grouped distributions, experiencing a notable decline in performance by 20.3% and 22.3%. This observed decrease in performance is consistent across models, unaffected by the specific positions of the labels within the prompt.
@@ 5 Conclusion
In summary, our research explores the capability of LLMs on long in-context learning tasks, particularly in extreme-label classification scenarios. We curate a dataset LongICLBench consisting of long in-context learning tasks with different difficulty levels in terms of context length. Through our study, we have discovered that LLMs demonstrate dramatic performance degradation when it comes to more difficult tasks. Our exploratory experiments further highlight the impact of the distribution of examples within prompts on model performance. We hope LongICLBench and our findings contribute to the ongoing efforts to enhance LLMs’ understanding of long contexts.
@@ A.2 Prompting Template
{comment: "…comment…"
emotion category: "…emotion…"
@@
{service query: "…service…"
intent category: "…intent…"
@@
{sentence: "…sentence…
the subject is "…subject…"
the object is "…object…"
the relation between the two entities is: "…relation…"
@@
{"…entity…": "…entity type…"
@@
"…dialogue…"
The list of entity pairs are "…(subject1, object1), (subject2, object2), etc…
The "…number of pairs…" respective relations between each entity pair are: "…relation, relation2, etc…
@@
{"…sentence1…" ( ) "…sentence2…"
the conjunction word in ( ) is "…conjunction…"
@@
A.3 Additional Distribution Analysis
To facilitate a clear comparison between random and grouped distributions, we organize instances of the same class to be adjacent within the demonstration prompts. The impact of this reorganization on model performance, both pre and post-grouping, is presented in Table 10.
Model Param Support Scatter Grouped $\Delta$
Context Tokens 10K
Gemma-7B-base 7B 8K 0 0 0
LLaMA-2-7B-32K 7B 32K 0.4 3.0 +2.6
ChatGLM3-6B-32K 6B 32K 38.9 35.6 -3.3
Qwen-1.5-7B-base 7B 32K 45.2 33.0 -12.2
Mistral-7B-v0.2-base 7B 32K 51.6 5.1 -46.5
LLaMA-2-7B-LongLora 7B 100K 0 0 0
Yi-6B-200K 6B 200K 8.0 0 -8
InternLM2-7B-base 7B 200K 15.5 4.8 -9.7
Long-LLaMA-code-7B 7B 256K 4.1 0 -4.1
RWKV-5-World 7B 4K 1.0 3.6 +2.6
Mamba-2.8B 2.8B 2K 0 0 0
GPT4-turbo N/A 128K 79.5 59.2 -20.3
Gemini-1.5-Pro N/A 10M 79.6 57.3 -22.3
The distribution plots for other models are shown in Figure 5 and Figure 6.
A.4 Data Accessibility
Our LongICLBench is set under MIT license, thus permission is granted, free of charge, to any person obtaining a copy of this dataset and associated documentation files. The datasests are curated under the rules guaranteed by the original dataset. There is no personally identifiable or offensive content in the dataset.
A.5 Broader Society Impacts
The development of long LLMs evaluation benchmarks and the corresponding insights can boost the development of long-context techniques, which can revolutionize fields requiring deep contextual understanding, such as legal analysis, long-form journalistic content generation, and comprehensive academic summarization. However, there are potential risks associated with the deployment of such powerful models. Enhanced long-context capabilities could be misused for generating misinformation, especially in political or social contexts, where nuanced long-form content can have significant influence. There is also the risk of dependency on automated systems in critical decision-making processes, which could lead to over-reliance on technology at the expense of human judgment.
A.6 Limitation
While the proposed LongICLBench can reveal the fact that current long LLMs still struggle with long-context understanding, it has limitations that LongICLBench currently encompasses only one type of evaluation application, specifically on extreme-label classification with long in-context learning. The intricacies involved in other novel long-context tasks that require a full understanding of the whole long-context are waiting to be put forward.